You have $25 to spend on picking 21 pounds of three different types of apples in an orchard. The Empire Apples cost $1.40 per pound, the Red Delicious Apples cost $1.10 per pound, and the Golden Delicious Apples cost $1.30 per pound. You want twice as many Red Delicious as the other two types combined. Write and solve a system of three equations to determine how many pounds of each type of apple you should buy
Answer by scott8148(6628)

let e=empire, r=red, and g=golden ... e+r+g=21 ... 14e+11r+13g=250 ... r=2(e+g)